# Exploring The Missing Semantics In Event Modality

### v1

@@ IV-C3 Qualitative Analysis on Semantic Feature Learning
IV-D Ablation Study
In this section, we provide a comprehensive quantitative evaluation of Semantic-E2VID, focusing on both the effectiveness of its designed components and its robustness under challenging conditions. We begin with detailed ablation studies to analyze the contribution of each core module—Learning Semantics, Fusing Semantics, and Supervising Semantics—by systematically replacing or modifying them and measuring the resulting performance changes. We further extend the analysis to key hyperparameters, such as the loss weighting factor $\lambda$ and the top-$N$ semantic categories used in supervision, to verify that our method remains stable across a range of parameter settings. Finally, we assess the robustness of Semantic-E2VID against real-world challenges, including varying event sparsity, different frame reconstruction intervals, and increased temporal irregularity. Results on the ECD and MVSEC datasets consistently show that Semantic-E2VID achieves the lowest MSE and highest SSIM compared to state-of-the-art baselines, highlighting its ability to produce accurate, stable, and semantically consistent reconstructions even under adverse conditions.
TABLE III: Results on ECD, MVSEC, and HQF datasets with different ablation settings.
@@ IV-D Ablation Study
Learning Semantics Direct Distillation 0.028 0.569 0.223 0.087 0.302 0.501 0.038 0.524 0.234
@@
Fusing Semantics Addition 0.025 0.575 0.220 0.078 0.305 0.500 0.036 0.535 0.228
@@
Supervising Semantics Original LPIPS 0.020 0.582 0.227 0.072 0.313 0.505 0.033 0.539 0.245
@@
IV-D1 Ablation Study on Designed Modules
To comprehensively assess the effectiveness of each proposed component—namely the Cross-modal Feature Alignment (CFA) module, the Semantic-guided Feature Fusion (SFF) module, and the Semantic Perceptual E2V Supervision—we perform a series of ablation studies, as summarized in Table III. The Full Model integrates all three components.
Learning Semantics. This ablation investigates the impact of the CFA module. In the comparison setting, Direct Distillation directly transfers the SAM-derived features $\mathcal{F}_{sam}$ into the event feature $\mathcal{F}_{e}$ output by the event encoder via relational distillation [41], without leveraging the CFA module. The Semantic Perceptual E2V Supervision remains enabled. While this direct distillation superficially incorporates semantic knowledge, it introduces two critical issues. First, $\mathcal{F}_{sam}$ originates from the RGB frame domain, whereas $\mathcal{F}_{e}$ captures high-dimensional features from asynchronous event data. The naive injection of frame-based semantics into the event feature space causes modality misalignment, which misleads the event encoder during training. Consequently, the event encoder learns distorted or semantically inconsistent representations, harming the reconstruction fidelity.
@@ IV-D1 Ablation Study on Designed Modules
Fusing Semantics. This study evaluates different strategies for integrating semantically enriched event features into the original event representations, while disabling the SFF module. Three fusion methods are considered:
@@
Supervising Semantics. This ablation examines the role of the Semantic Perceptual E2V Supervision. Here, we replace the proposed semantically guided loss with a standard perceptual loss, while keeping both the CFA and SFF modules intact. This setup allows us to assess the supervision signal’s contribution to guiding semantic reconstruction.
@@
IV-D2 Ablation Study on Hyperparameter
@@
Ablation on Hyperparameter in Loss Function. We conduct a robustness study on the hyper-parameter $\lambda$ in the final loss function (Eq. 6). Table IV reports the reconstruction performance on the ECD dataset across a range of $\lambda$ values. Our chosen value of $\lambda=1.8$ achieves the best performance in all three metrics—MSE, SSIM, and LPIPS. Notably, the performance variation is small across nearby values, demonstrating that our loss function is not overly sensitive to the exact choice of $\lambda$, which highlights the robustness of our formulation.
TABLE IV: Effect of hyper-parameter $\lambda$ in the final loss function $\mathcal{L}$. Results are reported on the ECD dataset.
@@
Ablation on Hyperparameter for Number of Masks from SAM. To evaluate the impact of the number of semantic mask categories used in our semantic perceptual E2V supervision, we conduct ablation experiments by varying the top-$N$ non-background semantic classes extracted from SAM. Table V reports the performance with $N\in\{5,7,10,15,20\}$ on the ECD dataset. Using too few semantic masks (e.g., $N=5$) fails to capture rich object-level guidance, leading to degraded reconstruction. On the other hand, using too many masks (e.g., $N=20$) introduces noisy or redundant categories that may weaken the supervision signal. Our default choice $N=10$ achieves the best balance across all metrics (MSE, SSIM, and LPIPS), validating its effectiveness in guiding the E2V reconstruction with semantically meaningful perceptual cues.
TABLE V: Ablation on the top-$N$ mask categories in semantic perceptual E2V supervision. Results are reported on the ECD dataset.
@@ IV-D3 Robustness Analysis
V Limitations and Future Work
@@
VI Conclusion
This paper presents Semantic-E2VID, a semantic-aware framework for event-to-video reconstruction. Leveraging the semantic knowledge from a visual foundation model (SAM), our method bridges the gap in event modality lacking semantic information and learns high-level semantic representations from sparse event data. The Cross-Modal Feature Alignment (CFA) module extracts rich semantics, while the Semantic-aware Feature Fusion (SFF) block integrates them with event features for enhanced reconstruction. A Semantic Perceptual E2V Supervision further guides semantic-level accuracy. Experiments show that Semantic-E2VID achieves superior results across diverse scenarios, particularly under extreme lighting. This work demonstrates the benefit of incorporating external semantic priors into event-based vision and points toward future exploration of event-based visual foundation models.

### v2

@@ 4.3.4 Qualitative Analysis on Semantic Feature Learning
4.4 Ablation Study
In this section, we provide a comprehensive quantitative evaluation of Semantic-E2VID, focusing on both the effectiveness of its designed components and its robustness under challenging conditions. We begin with detailed ablation studies to analyze the contribution of each core module—Semantic Abstraction, Semantic Feature Fusing, and Semantic-Aware Supervision-replacing or modifying them and measuring the resulting performance changes. We further extend the analysis to key hyperparameters, such as the loss weighting factor $\lambda$ and the top-$N$ semantic categories used in supervision, to verify that our method remains stable across a range of parameter settings. Finally, we assess the robustness of Semantic-E2VID against real-world challenges, including varying event sparsity, different frame reconstruction intervals, and increased temporal irregularity. Results on the ECD and MVSEC datasets consistently show that Semantic-E2VID achieves the lowest MSE and highest SSIM compared to state-of-the-art baselines, highlighting its ability to produce accurate, stable, and semantically consistent reconstructions even under adverse conditions.
Table 3: Results on ECD, MVSEC, and HQF datasets with different ablation settings.
@@ 4.4 Ablation Study
Semantic Abstraction Direct Distillation 0.028 0.569 0.223 0.087 0.302 0.501 0.038 0.524 0.234
@@
Semantic Feature Fusing Addition 0.025 0.575 0.220 0.078 0.305 0.500 0.036 0.535 0.228
@@
Semantic-Aware Supervision Original LPIPS 0.020 0.582 0.227 0.072 0.313 0.505 0.033 0.539 0.245
@@
4.4.1 Ablation Study on Designed Modules
To comprehensively assess the effectiveness of each proposed component—namely the Cross-modal Feature Alignment (CFA) module, the Semantic-guided Feature Fusion (SFF) module, and the Semantic Perceptual E2V Supervision—we perform a series of ablation studies, as summarized in Table 3. The Full Model integrates all three components.
Semantic Abstraction. This ablation investigates the impact of the CFA module. In the comparison setting, Direct Distillation directly transfers the SAM-derived features $\mathcal{F}_{sam}$ into the event feature $\mathcal{F}_{e}$ output by the event encoder via relational distillation [29], without leveraging the CFA module. The Semantic Perceptual E2V Supervision remains enabled. While this direct distillation superficially incorporates semantic knowledge, it introduces two critical issues. First, $\mathcal{F}_{sam}$ originates from the RGB frame domain, whereas $\mathcal{F}_{e}$ captures high-dimensional features from asynchronous event data. The naive injection of frame-based semantics into the event feature space causes modality misalignment, which misleads the event encoder during training. Consequently, the event encoder learns distorted or semantically inconsistent representations, harming the reconstruction fidelity.
@@ 4.4.1 Ablation Study on Designed Modules
Semantic Feature Fusing. This study evaluates different strategies for integrating semantically enriched event features into the original event representations, while disabling the SFF module. Three fusion methods are considered:
  • 1.
@@
Semantic-Aware Supervision. This ablation examines the role of the Semantic Perceptual E2V Supervision. Here, we replace the proposed semantically guided loss with a standard perceptual loss, while keeping both the CFA and SFF modules intact. This setup allows us to assess the supervision signal’s contribution to guiding semantic reconstruction.
@@
4.4.2 Ablation Study on Hyperparameter
@@
Ablation on Hyperparameter in Loss Function. We conduct a robustness study on the hyper-parameter $\lambda$ in the final loss function (Eq. 6). Table 4 reports the reconstruction performance on the ECD dataset across a range of $\lambda$ values. Our chosen value of $\lambda=1.8$ achieves the best performance in all three metrics—MSE, SSIM, and LPIPS. Notably, the performance variation is small across nearby values, demonstrating that our loss function is not overly sensitive to the exact choice of $\lambda$, which highlights the robustness of our formulation.
Table 4: Effect of hyper-parameter $\lambda$ in the final loss function $\mathcal{L}$. Results are reported on the ECD dataset.
@@
Ablation on Hyperparameter for Number of Masks from SAM. To evaluate the impact of the number of semantic mask categories used in our semantic perceptual E2V supervision, we conduct ablation experiments by varying the top-$N$ non-background semantic classes extracted from SAM. Table 5 reports the performance with $N\in\{5,7,10,15,20\}$ on the ECD dataset. Using too few semantic masks (e.g., $N=5$) fails to capture rich object-level guidance, leading to degraded reconstruction. On the other hand, using too many masks (e.g., $N=20$) introduces noisy or redundant categories that may weaken the supervision signal. Our default choice $N=10$ achieves the best balance across all metrics (MSE, SSIM, and LPIPS), validating its effectiveness in guiding the E2V reconstruction with semantically meaningful perceptual cues.
Table 5: Ablation on the top-$N$ mask categories in semantic perceptual E2V supervision. Results are reported on the ECD dataset.
@@ 4.4.3 Robustness Analysis
5 Limitations and Future Work
@@
6 Conclusion
In this work, we revisited event-to-video reconstruction from a semantic perspective and identified semantic under-determinacy as a fundamental limitation of event-based sensing. Unlike conventional E2V approaches that rely primarily on temporal aggregation and spatial feature modeling, we argued that faithful reconstruction from sparse event streams requires explicitly accounting for missing object-level structure and contextual information. Based on this insight, we proposed Semantic-E2VID, a semantic-enriched end-to-end E2V framework that reformulates reconstruction as a process of semantic abstraction, semantic feature fusion, and semantic-aware supervision. By bridging event representations with semantic knowledge extracted from a pretrained foundation model in a representation-compatible manner, our approach enables event features to capture object-level structure and contextual cues that are missing in raw event streams. Extensive experiments on six public benchmarks demonstrate that Semantic-E2VID consistently outperforms state-of-the-art methods, producing reconstructions with clearer object boundaries, stronger semantic consistency, and improved robustness under sparse or ambiguous event conditions.